a company makes a range of products with total budgeted manufacturing overheads of $900,000 incurred in three production (a,b, and c) and one service department.

department a has 11 direct employees who each work 35 hours per week. department b has 7 machines each of which is operated for 24 hours per week. department c is expected to produce 148,000 units of final product in the budget period.the company will operate for 50 weeks in the budget period.budgeted overheads incurred directly by each department are:

production department a $261,745

department b $226,120

department c $93,890

service department d $53,305

the balance of budgeted overheads are apportioned to departments as follows:

production department a 40%

production department b 35%

production department c 20%

service department d 5%

service department overheads are apportioned equally to each production department require:

(a)calculate an appropriate predetermined overhead absorption rate in each production department.

(b)calculate the manufacturing overhead cost per unit of finished product in batch of 100 units which take nine direct labour hours in department a and three machine hours in department b to produce.,
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
A company
Calculation of overheads to be apportioned: Amount $ Note
Total budgeted manufacturing overhead 900,000.00
Less: incurred directly by each department
Production department a 261,745.00
Production department b 226,120.00
Production department c      93,890.00
Service department      53,305.00
Overheads to be apportioned: 264,940.00 A
Apportionment of remaining overheads: Dept. a Dept. b Dept. c Service Total Note
Ratio 40% 35% 20% 5% 100% B
Overheads apportioned: 105,976.00      92,729.00      52,988.00      13,247.00 264,940.00 C=B*A
Apportionment of Service dept. overhead Amount $ Note
Overhead incurred directly      53,305.00
Overhead apportioned      13,247.00 See C
Total Overhead     66,552.00 D
Share of each production department     22,184.00 E=D/3
Total overhead of each production dept. Dept. a Dept. b Dept. c Total Note
Overhead incurred directly 261,745.00 226,120.00      93,890.00 581,755.00
Overhead apportioned 105,976.00      92,729.00      52,988.00 251,693.00 See C
Share of Service dept. overhead      22,184.00      22,184.00      22,184.00     66,552.00 See E
Total overhead of each production dept. 389,905.00 341,033.00 169,062.00 900,000.00 F
Total labor hours of dept. a. Note
Number of employees             11.00 G
Hours worked per week by each employee             35.00 H
Total labor hours worked per week           385.00 I=G*H
Number of weeks operated             50.00 J
Total labor hours worked     19,250.00 K=I*J
Answer a
Predetermined overhead absorption rate in department a             20.25 L=F/K
Total machine hours of dept. b. Note
Number of machines                7.00 M
Hours operated per week by each machine             24.00 N
Total hours operated in a week           168.00 O=M*N
Number of weeks operated             50.00 P
Total labor hours worked        8,400.00 Q=O*P
Answer a
Predetermined overhead absorption rate in department b             40.60 R=F/Q
Answer a
Number of units produced in dept. c. 148,000.00 S
Predetermined overhead absorption rate in department c                1.14 T=F/S
Answer b Dept. a Dept. b Dept. c Total Note
Activities used             11.00                3.00           100.00 U
Predetermined overhead absorption rate             20.25             40.60                1.14 See L, R, T
Allocated manufacturing overhead           222.80           121.80           114.23           458.83 This is Activities used* Predetermined overhead absorption rate.
Total units           100.00
Manufacturing overhead cost per unit of finished product                4.59 This is Allocated manufacturing overhead/ Total units.